### v1

@@ Abstract
Gender pay equity remains an open challenge in academia despite decades of movements. Prior studies, however, have relied largely on descriptive regressions, leaving causal analysis underexplored. This study examines gender-based wage disparities among tenure-track faculty in the University of North Carolina system using both parametric and non-parametric causal inference methods. In particular, we employed propensity score matching and causal forests to estimate the causal effect of gender on academic salary while controlling for university type, discipline, titles, working years, and scholarly productivity metrics. The results indicate that on average female professors earn approximately 6% less than their male colleagues with similar qualifications and positions.
Keywords: Gender wage gap, Higher education, Causal inference, Academic labor markets
@@ 1 Introduction
Understanding the gender wage gap has long been a central question in labor economics and the social sciences. Since the foundational work of Blinder (1973) and Oaxaca (1973), researchers have debated whether wage disparities reflect differences in productivity or are rooted in systemic bias. The academic labor market, which prides itself on meritocracy and performance-based advancement, presents a particularly puzzling case: substantial evidence shows that women faculty continue to earn less than their male counterparts, even after accounting for key observable factors such as rank, field, and scholarly output (Ginther and Kahn, 2004; Ceci et al., 2014; Mishel and Schieder, 2016; Monroe et al., 2008). These gaps persist despite the growing presence of women in academia and widespread institutional efforts to promote equity in hiring and advancement (Blau and Kahn, 2017; U.S. Bureau of Labor Statistics, 2023).
@@
Two main perspectives have shaped how scholars understand gender pay gaps in academia. One line of thinking points to structural factors: women are more likely to hold lower-paid positions—such as assistant professorships, non-tenure-track jobs, or teaching-intensive roles—which naturally brings down their average earnings (Ginther and Kahn, 2004; Monroe et al., 2008). Another view focuses on differences in human capital, suggesting that salary disparities reflect variations in research productivity, often measured through publication counts or citation impact (Ceci et al., 2014; Kim et al., 2024). While some studies—such as Ceci et al. (2014)—find relatively small pay differences after accounting for productivity, their work still treats scholarly output as the core metric of academic value. In practice, however, it can be difficult to cleanly separate these two explanations, since rank, research output, and compensation tend to evolve together throughout a faculty member’s career.
@@
Much of the existing empirical work relies on associational methods, such as Oaxaca–Blinder decompositions or pooled regressions using cross-sectional data. These approaches are limited in that they cannot estimate counterfactual salary outcomes, nor can they fully account for unobserved confounding—such as career interruptions or uneven service burdens—that may influence both productivity and compensation.
@@
In this study, we address these limitations by applying modern causal inference tools to the gender wage gap among professors affiliated to institutions in the North Carolina public university system. We merge faculty salary records from the University of North Carolina system, where all sixteen campuses are required by law to disclose annual salaries, to academic performance data from Google Scholar and gender information from genderize.io. Our dataset uses salary information from 2022 and includes 12,039 entries with detailed information on academic rank, department, institutional affiliation, working years, and research productivity, as measured by the log-transformed i10-index. This allows us to move beyond descriptive associations and estimate the causal effect of gender on log-transformed salary outcomes while adjusting for key covariates such as academic title, department, institutional affiliation, working years, and research productivity (measured via log-transformed i10-index). We also explore treatment effect heterogeneity by examining how the gender wage gap varies across career stages and research performance tiers.
By analyzing a rich dataset of publicly available faculty salaries and bibliometric indicators, our study contributes to ongoing discussions in several ways. First, we introduce a counterfactual framework to estimate the average and subgroup-specific effects of gender on salary. Second, we incorporate machine learning methods, such as causal forests, to uncover patterns of heterogeneous treatment effects that traditional parametric models may overlook. Our results not only quantify the magnitude of the gender wage gap, but also illuminates to its complex and context-dependent mechanisms. Identifying where disparities emerge and where they persist even after controlling for academic performance, is a necessary step toward realizing true pay equity in academia.
This article follows this structure: section two introduces the dataset and describes key variables; section three outlines the methodological framework, beginning with preliminary analyses and then detailing both parametric and non-parametric approaches to causal inference; section four presents the results, first offering descriptive and non-causal findings, followed by estimates from parametric and non-parametric causal models; section five concludes with a discussion of the main findings, limitations, and directions for future research.
@@ 2 Data and Variables
To investigate the relationship between gender and faculty salary, we compiled a large database of employee salaries from 2022 of all sixteen institutions in North Carolina public university system. These sixteen universities within the one university system share the same governance structure and operate under the same policy orientation including faculty hiring and promotion (The University of North Carolina, 2023). The data used in this study was obtained from the University of North Carolina Digital Measures website (The University of North Carolina, 2022). The dataset contains 47,405 entries, with information on full names, age, initial hired date, rank, department, and the salaries of all employees, from faculty members to administrative and support staff, affiliated with one of the institutions. We removed non-faculty employees (such as department staff, facilities employees, athletic coaches, etc.) from the data based on the employee home department and a column that indicates job functions. Also, the professors with less than $27,000$ yearly salary were eliminated (Lu and Hannig, 2024) as these are likely clerical errors. The resulting dataset contains 12,039 entries, focusing on tenure track faculty (assistant professor, associate professor, and full professor) in all institutions from North Carolina public university system.
Our study focuses on examining how gender influences the faculty salary. We constructed our dependent variable using annual faculty salary data. This variable was directly generated from the raw dataset. In our cleaned dataset, the average yearly salary is \$128,000. We use log10 transformation for our analysis because salary distributions tend to be right-skewed, see Figure 2, and the logarithmic transformation normalizes the data and reduces the influence of extreme values, Figure 2. Additionally, this transformation allows us to interpret the results in terms of percentage change rather than absolute dollar amounts, which is more meaningful when analyzing salary differentials across different faculty ranks and disciplines. This dependent variable (totally 12,039 entries) contains no missing values.
@@
Since our raw dataset does not contain faculty gender information, we followed established methodologies from the literature to infer gender based on first names (Kim et al., 2022). We employed the Genderize API  [^1] https://genderize.io/ to assign gender labels to our dataset. The API returns four key variables for each query: the predicted gender classification, the confidence score of the prediction ranging from 0–100%, the number of samples used for the prediction, and the queried first name. To ensure high accuracy in our gender assignments, we implemented a two-stage process. First, we used the Genderize API to assign gender automatically for cases with confidence scores greater than or equal to 60%. Second, we performed manual verification for the remaining cases, including names with confidence scores below 60%, cases where gender remains undetermined, and uncommon or culturally specific names. The number of cases dealt with each methods are shown in Table A.1. This hybrid approach, combining algorithmic prediction with manual verification, follows best practices established in recent literature (Wu and Smith, 2020; Karimi and Wang, 2016) and helps reduce potential biases in automated gender inference systems. The manual verification process involved considering faculty websites, professional profiles, and institutional directories to determine gender with certainty. After this hybrid solution, there is no missing data for gender. In this dataset, 45.5% of the entries correspond to females, while 54.5% represent males.
Naturally, there are many variables that influence academic salaries. These variables were treated as confounding factors in our analysis.
The first confounding variable we considered is length of employment. To consider career progression effects on salary, we calculated working years as the difference between 2022 and each faculty member’s initial hire date. The average working years in our sample is $14.6$ years (SD = $9.82$). The variable working years has no missing values.
@@
Then, we included title as another confounding variable, since title information helps control for career stage, which strongly influences salary levels (Toutkoushian et al., 2007). This information is provided in the raw dataset and no missing values. We removed non-faculty employees and only kept entries with titles containing professor. Then, we categorized faculty titles into three groups: Assistant Professor (35.1%), Associate Professor (31.0%), and Professor (34.0%).
@@
To account for effects related to the various missions of the 16 universities within the UNC system, we used variable university classification. Following Carnegie Classification of Institutions of Higher Education [^2] https://carnegieclassifications.acenet.edu/, which is a widely used framework for categorizing colleges and universities in the United States, we categorized these institutions into four types: bachelor (3 institutions), master (5 institutions), DRU(H) (6 institutions), and DU/VA (2 institutions). The first category is Bachelor’s Colleges and Universities (bachelor), in which the institutions award bachelor’s degrees to at least 50% of their graduates, while awarding fewer than 50 master’s degrees and fewer than 20 doctoral degrees annually. The second is Master’s Colleges and Universities (master). These institutions award at least 50 master’s degrees annually, while awarding fewer than 20 doctoral degrees annually. For Doctoral/Research Universities (DRU(H)), the universities award at least 20 research/scholarship doctoral degrees annually (excluding professional degrees such as MD, JD, PharmD). The last category is Doctoral Universities or Very High Research Activity (DU/VA), where the institutions are characterized by very high levels of research activity and a significant output of doctoral degrees in diverse fields, often referred to as R1 universities. For our analysis, we combined bachelor’s and master’s institutions into a single category, as both types of institutions share similar characteristics in terms of their primary focus on teaching rather than research (Henderson and Buchanan, 2007). After assigning each university to its according university code, we got a more comprehensive dataset without any missing value in this new column. These classifications also reflect the distribution of faculty members across NC public university system: 47.4% are affiliated with R1 universities (DU/VA), 38.6% with R2 universities (DRU(H)), and 14.0% with primarily undergraduate or master’s institutions (bachelor/master).
@@
Controlling for faculty’s home department is crucial because academic disciplines significantly influence salary levels, with substantial variations observed across fields (Ehrenberg, 2004). We categorized departments into six broad fields: Business (6.5%), Technology and Engineering (15.3%), Arts and Humanities (18.1%), Medical and Health Sciences (27.2%), Natural Sciences (16.3%), and Social Sciences (16.6%). This categorization follows the UNESCO’s International Standard Classification of Education (ISCED-F) framework (UNESCO Institute for Statistics, 2014), which provides a standardized system for classifying academic disciplines in higher education. We used Large Language Models (LLMs), specifically ChatGPT, for department name standardization. LLMs, based on transformer architectures (Vaswani et al., 2017), have demonstrated superior performance in understanding semantic relationships and contextual variations in text (Brown et al., 2020). This approach enabled efficient and accurate mapping of diverse department names into our predefined categories while maintaining consistency across institutions.
@@
As academic performance may also significantly influence faculty compensation (Fairweather, 2005). To account for this we matched faculty members to their Google Scholar profiles when available (match rate = 41.4%) by using Google Scholar API. This identifier enables us to collect publication metrics and establish research productivity measures. From Google Scholar, we extracted several metrics for each faculty member: total citations (citedby), 5-year citations (citedby5y), h-index (hindex), 5-year h-index (hindex5y), i10-index (i10index), and 5-year i10-index (i10index5y). To avoid missing values due to the max access limitations of Google Scholar API, we divided the data into smaller files combining them into one complete file at the end. Despite implementing this well-designed solution, a potential issue remains: the possibility of retrieving information for faculty members with identical names rather than the intended individual. To address this, we incorporated an additional validation step using the email domain provided by the Google Scholar API. Specifically, we matched the email domain with the university’s domain to ensure accuracy. Entries with matching domains were retained, while those without a match were excluded. However, this approach introduces a new limitation. Faculty members who have used an email address associated with a previous institution may be incorrectly excluded, as the domain would not align with their current university. To avoid the limitation caused by the email matching exclusion, we manually tested the excluded entries. So, the three-step testing method provides a more reliable linkage between faculty and their profiles, ensuring higher overall accuracy. Among academic metrics from Google Scholar API, we selected i10-index, which measures research impact by counting publications with at least 10 citations (Google Scholar, 2011), as our primary measure of research productivity because it captures both the quantity and impact of scholarly work. We also tried other academic indicators such as h-index but in our models athe i10-index was the best metric for academic performances. We applied log10 transformation to i10-index because citation metrics typically follow a highly skewed distribution (Evans and Steptoe-Warren, 2019), shown in Figure 4. The log transformation helps normalizing the distribution and reduces the influence of extreme values in our analysis, see Figure 4; the mean log10(i10-index) is 1.36 (SD = 0.456). For faculties without Google Scholar ID, we imputed their log10(i10index) using the mean value of similar background faculty members who had available academic information. Using this metric helps control for research productivity’s influence on salary, as higher research output and citation impact are associated with higher faculty compensation (Gibson et al., 2014).
@@ 3.2 Causal Inference Framework
Causal inference methods aim to estimate the effect of a treatment (in this case, gender) on an outcome (salary), while addressing confounding. In our context, we treated gender as a binary treatment variable ($Z_{i}$), where $Z_{i}=1$ for female and $Z_{i}=0$ for male, and $Y_{i}$ represents the observed salary. To adjust for potential confounding, we condition on a set of observed covariates $X_{i}$, which are associated with both gender and salary. In our study, $X_{i}$ consists of academic title, department code, university code, years of experience, log-transformed i10-index, and Google Scholar ID. These covariates capture institutional characteristics, professional experience, and research productivity.
@@
This formulation inherently relies on the concept of counterfactuals: for each individual, we observe only one of the two potential outcomes—either as male or female—but not both. The unobserved outcome represents the counterfactual, and causal inference techniques are designed to estimate this missing information using observed data and plausible assumptions (Imbens and Rubin, 2015).
@@
    Consistency and SUTVA: The observed outcome equals the potential outcome corresponding to the treatment received:
@@
    This also assumes that treatment is well-defined and that there is no interference between units (Stable Unit Treatment Value Assumption, or SUTVA).

### v2

@@ Abstract
Observational studies often present challenges for causal inference due to confounding and heterogeneity. In this paper, we illustrate how modern causal inference methods can be applied to large-scale academic salary data. Using records from 12,039 tenure-track faculty in the University of North Carolina system, linked with bibliometric indicators and institutional classifications, we estimate the causal effect of gender on faculty salaries. Our analysis combines propensity score matching with causal forests to adjust for rank, discipline, research productivity, and career experience. Results indicate that female faculty earn approximately 6% less than comparable male colleagues, with variation in the gap across career stages and levels of research productivity. This case study demonstrates how causal inference methods for observational data can provide insight into structural disparities in complex social systems.
Keywords: Causal inference, Observational data, Propensity score matching, Causal forests, Gender wage inequality, Higher education
@@ 1 Introduction
Understanding and explaining the structure and persistence of the gender wage gap remains a central question in labor economics and the social sciences. Since the foundational work of blinder1973wage and oaxaca1973male, the literature has sought to distinguish disparities attributable to productive characteristics from those arising through labor-market institutions. The academic labor market, which prides itself on meritocracy and performance-based advancement, presents a particularly puzzling case: despite controlling for rank, field, and research output, standard wage equations consistently reveal lower earnings for women faculty relative to their male counterparts (ginther2004women; ceci2014women; monroe2008gender), yet this gap has rarely been examined using causal estimators. These disparities persist despite the growing representation of women in academia and widespread institutional efforts to promote equity in hiring and advancement (blau2017gender; BLS2023).
@@
Two main perspectives have shaped scholarly understanding of gender pay gaps in academia. One attributes disparities to structural factors: women are more likely to occupy lower-paid positions—such as assistant professorships, non-tenure-track roles, or teaching-intensive appointments—which naturally depresses their average earnings (ginther2004women; monroe2008gender). Another perspective emphasizes differences in human capital, suggesting that salary disparities reflect variation in research productivity, often measured through publication counts or citation impact (ceci2014women; kim2024gendergap). While some studies—such as ceci2014women—report relatively small pay differences after accounting for productivity, their work continues to treat scholarly output as the primary metric of academic value. In practice, however, it is difficult to disentangle these explanations, as rank, research output, and compensation typically evolve in tandem over the course of a faculty member’s career.
@@
Much of the existing empirical literature relies on associational methods, such as Oaxaca–Blinder decompositions or pooled regressions based on cross-sectional data. These approaches are limited in that they cannot estimate counterfactual salary outcomes, nor can they fully account for unobserved confounding—such as career interruptions or uneven service burdens—that may influence both productivity and compensation. While prior work has documented the existence of gender pay disparities in academia, there remains limited causal evidence on how these gaps vary by institutional type, career stage, or research performance. These gaps are central to identifying equity interventions.
@@
In this study, we address these limitations by applying modern causal inference tools to examine the gender wage gap among professors in the North Carolina public university system. We merge faculty salary records from all sixteen campuses in the University of North Carolina system—which are legally required to disclose annual salaries—with academic performance data from Google Scholar and gender labels inferred via genderize.io. Our dataset covers salary information from 2022 and includes 12,039 faculty, with detailed information on academic rank, department, institutional affiliation, years of experience, and research productivity, measured using the log-transformed i10-index. This rich dataset enables us to move beyond descriptive associations and address core challenges in observational causal inference, such as confounding, covariate imbalance, and treatment effect heterogeneity, by estimating the causal effect of gender on log-transformed salary outcomes. We also explore treatment effect heterogeneity by examining how the gender wage gap varies across career stages and levels of research performance.
Our study contributes to the literature on gender inequality and academic labor markets in three key ways. First, we provide a causal estimate of the gender wage gap across a comprehensive public university system, offering a more generalizable view of pay disparities than analyses limited to a single institution or field. Second, we move beyond average treatment effects by applying causal forests to uncover patterns of heterogeneity in the gender wage gap. This method reveals how the magnitude of pay disparities varies across career stages and levels of research productivity, offering new insights into where interventions may be most needed. Third, we release a reproducible workflow that integrates public salary disclosures, bibliometric data from Google Scholar, and modern causal inference techniques. This framework can be readily adapted for equity audits in other academic systems.
The remainder of the paper proceeds as follows. Section 2 describes the dataset and variable construction. Section 3 outlines the methodological framework, including both parametric and non-parametric approaches. Section 4 presents the empirical findings. Section 5 concludes with implications and future directions.
@@ 2 Data and Variables
To investigate the relationship between gender and faculty salary, we compiled a large database of employee salaries from 2022 of all sixteen institutions in North Carolina public university system. These sixteen universities within the one university system share the same governance structure and operate under the same policy orientation including faculty hiring and promotion (UNCpolicy2023). The data used in this study was obtained from the University of North Carolina Digital Measures website (UNCsalary2022). The dataset contains 47,405 entries, with information on full names, age, initial hired date, rank, department, and the salaries of all employees, from faculty members to administrative and support staff, affiliated with one of the institutions. We removed non-faculty employees (such as department staff, facilities employees, athletic coaches, etc.) from the data based on the employee home department and a column that indicates job functions. Also, the professors with less than $27,000$ yearly salary were eliminated (lu2024gender) as these are likely clerical errors. In total, seven observations were dropped from the dataset. The resulting dataset contains 12,039 entries, focusing on tenure track faculty (assistant professor, associate professor, and full professor) in all institutions from North Carolina public university system.
Our study focuses on examining how gender influences the faculty salary. We constructed our dependent variable using annual faculty salary data. This variable was directly generated from the raw dataset. In our cleaned dataset, the average yearly salary is \$128,000. We use a log10 transformation for our analysis because salary distributions tend to be right-skewed, see Figure 2, and the logarithmic transformation normalizes the data and reduces the influence of extreme values, Figure 2. Additionally, this transformation allows us to interpret the results in terms of percentage change rather than absolute dollar amounts, which is more meaningful when analyzing salary differentials across different faculty ranks and disciplines. Importantly, faculty appointments in the North Carolina public university system are explicitly categorized into either nine- or twelve-month “service periods” (unc_facultyaffairs_service_periods). Most teaching and research faculties hold standard nine-month academic-year contracts, making their base salaries directly comparable. For the relatively few twelve-month appointments (e.g., in medical schools), any salary differences reflect discipline-specific norms and are absorbed by our department fixed effects. This dependent variable (totally 12,039 entries) contains no missing values.
@@
Since our raw dataset does not contain faculty gender information, we followed established methodologies from the literature to infer gender based on first names (kim2022gendered). We employed the Genderize API  [^1] https://genderize.io/ to assign gender labels to our dataset. The API returns four key variables for each query: the predicted gender classification, the confidence score of the prediction ranging from 0–100%, the number of samples used for the prediction, and the queried first name. To ensure high accuracy in our gender assignments, we implemented a two-stage process. First, we used the Genderize API to assign gender automatically for cases with confidence scores greater than or equal to 60%. Second, we performed manual verification for the remaining cases, including names with confidence scores below 60%, cases where gender remains undetermined, and uncommon or culturally specific names. The number of cases dealt with each methods are shown in Table A.1. This hybrid approach, combining algorithmic prediction with manual verification, follows best practices established in recent literature (vanhelene2024inferring) and helps reduce potential biases in automated gender inference systems. The manual verification process involved considering faculty websites, professional profiles, and institutional directories to determine gender with certainty. After this hybrid solution, there is no missing data for gender. In this dataset, 45.5% of the entries correspond to females, while 54.5% represent males.
@@
Naturally, there are many variables that influence academic salaries. These variables were treated as confounding factors in our analysis. The first confounding variable we considered is length of employment. To consider career progression effects on salary, we calculated working years as the difference between 2022 and each faculty member’s initial hire date. The average working years in our sample is $14.6$ years (SD = $9.82$). The variable working years has no missing values. Next, we included title as another confounding variable, since title information helps control for career stage, which strongly influences salary levels (griswold1996political). This information is provided in the raw dataset and has no missing values. We removed non-faculty and non-tenure track employees and categorized faculty titles into three groups: Assistant Professor (35.1%), Associate Professor (31.0%), and Professor (34.0%).
@@
To account for effects related to the various missions of the 16 universities within the UNC system, we used variable university classification. Following Carnegie Classification of Institutions of Higher Education [^2] https://carnegieclassifications.acenet.edu/, which is a widely used framework for categorizing colleges and universities in the United States, we categorized these institutions into four types: bachelor (3 institutions), master (5 institutions), DRU(H) (6 institutions), and DU/VA (2 institutions). The first category is Bachelor’s Colleges and Universities (bachelor), in which the institutions award bachelor’s degrees to at least 50% of their graduates, while awarding fewer than 50 master’s degrees and fewer than 20 doctoral degrees annually. The second is Master’s Colleges and Universities (master). These institutions award at least 50 master’s degrees annually, while awarding fewer than 20 doctoral degrees annually. For Doctoral/Research Universities (DRU(H)), the universities award at least 20 research/scholarship doctoral degrees annually (excluding professional degrees such as MD, JD, PharmD). The last category is Doctoral Universities or Very High Research Activity (DU/VA), where the institutions are characterized by very high levels of research activity and a significant output of doctoral degrees in diverse fields, often referred to as R1 universities. For our analysis, we combined bachelor’s and master’s institutions into a single category, as both types of institutions share similar characteristics in terms of their primary focus on teaching rather than research (henderson2007scholarship). Across the UNC system: 47.4% are affiliated with R1 universities (DU/VA), 38.6% with R2 universities (DRU(H)), and 14.0% with primarily undergraduate or master’s institutions (bachelor/master).
@@
Controlling for faculty’s home department is crucial because academic disciplines significantly influence salary levels, with substantial variations observed across fields (toutkoushian2007interaction). We categorized departments into six broad fields: Business (6.5%), Technology and Engineering (15.3%), Arts and Humanities (18.1%), Medical and Health Sciences (27.2%), Natural Sciences (16.3%), and Social Sciences (16.6%). This categorization follows the UNESCO’s International Standard Classification of Education (ISCED-F) framework (unesco2014isced), which provides a standardized system for classifying academic disciplines in higher education. We used Large Language Models (LLMs), specifically ChatGPT, for department name standardization. LLMs, based on transformer architectures (vaswani2017attention), have demonstrated superior performance in understanding semantic relationships and contextual variations in text (brown2020language). This approach enabled efficient and accurate mapping of diverse department names into our predefined categories while maintaining consistency across institutions. To ensure reliability, a random sample of ChatGPT-generated classifications was manually reviewed and confirmed.
@@
As academic performance may also significantly influence faculty compensation (fairweather2005beyond). To account for this we matched faculty members to their Google Scholar profiles when available (match rate = 41.4%) by using Google Scholar API. This identifier enables us to collect publication metrics and establish research productivity measures. From Google Scholar, we extracted several metrics for each faculty member: total citations (citedby), 5-year citations (citedby5y), h-index (hindex), 5-year h-index (hindex5y), i10-index (i10index), and 5-year i10-index (i10index5y). To deal with the max access limitations of Google Scholar API, we divided the data into smaller files combining them into one complete file at the end. In order to address the possibility of retrieving information for another faculty member with identical names rather than the intended individual, we incorporated an additional validation step using the email domain provided by the Google Scholar API. Specifically, we matched the email domain with the university’s domain to ensure accuracy. Entries with matching domains were retained, while those without a match were excluded. However, this approach introduces a new limitation. Faculty members who have used an email address associated with a previous institution may be incorrectly excluded, as the domain would not align with their current university. To avoid the limitation caused by the email matching exclusion, we manually tested the excluded entries. Consequently, the three-step testing method provides a more reliable linkage between faculty and their profiles, ensuring higher overall accuracy. Among academic metrics from Google Scholar API, we selected i10-index, which measures research impact by counting publications with at least 10 citations (google2011metrics), as our primary measure of research productivity because it captures both the quantity and impact of scholarly work. We also tried other academic indicators such as h-index but in our models the i10-index was the best metric for academic performance. We applied log10 transformation to i10-index because citation metrics typically follow a highly right-skewed distribution (hammouri2020skewedata), shown in Figure 4. The log transformation helps normalizing the distribution and reduces the influence of extreme values in our analysis, see Figure 4; the mean log10(i10-index) is 1.36 (SD = 0.456). For faculties without Google Scholar ID, we imputed their log10(i10index) using the mean value of similar background faculty members who had available academic information. Using this metric helps control for research productivity’s influence on salary, as higher research output and citation impact are associated with higher faculty compensation (wikipediaHindex).
@@ 3.2 Causal Inference Framework
Causal inference methods aim to estimate the effect of a treatment (in this case, gender) on an outcome (salary), while addressing confounding. In our context, we treated gender as a binary treatment variable ($Z_{i}$), where $Z_{i}=1$ for female and $Z_{i}=0$ for male, and $Y_{i}$ represents the observed salary. To adjust for potential confounding, we condition on a set of observed covariates $X_{i}$, which are associated with both gender and salary. In our study, $X_{i}$ consists of academic title, department code, university code, years of experience, log-transformed i10-index, and Google Scholar ID. These covariates capture institutional characteristics, professional experience, and research productivity. While gender is not a manipulable policy intervention, we follow the causal inference literature in treating it as an exposure variable within the potential outcomes framework. Under the assumption of conditional ignorability, our estimates can be interpreted as the causal effect of gender on salary outcomes (sloczynski2020).
@@
This formulation inherently relies on the concept of counterfactuals: for each individual, we observe only one of the two potential outcomes—either as male or female—but not both. The unobserved outcome represents the counterfactual, and causal inference techniques are designed to estimate this missing information using observed data and plausible assumptions (imbens2015causal).
@@
    Stable Unit Treatment Value Assumption (SUTVA): The observed outcome equals the potential outcome corresponding to the treatment received:
@@
    This also assumes that treatment is well-defined and that there is no interference between units.